OpenStudy (kaederfds):

If A+B=5 then the value of B-3+A is ...

OpenStudy (callisto):

B-3+A = A+B -3 = 5-3 =?

OpenStudy (kaederfds):

2!

OpenStudy (callisto):

You've found it :)